Required information

[The following information applies to the questions displayed below.]

On April 30, 2017, Tilton Products purchased machinery for $55,000. The useful life of this machinery is estimated at 8 years, with an $5,000 residual value.

1A. In the year 2023, Tilton Products sells this machinery for $4,000. At the date of sale, the machinery had been depreciated by Tilton Products to its estimated residual value of $5,000. This sale results in:

1B. Assume that in its financial statements, Tilton Products uses the 200%-declining-balance method and the half-year convention. Depreciation expense in 2017 and 2018 will be:
Answer #1

1A.
Since book value of machinery at time of sale = $5000, sale proceeds = $4000
There will be loss on sale of machinery = $5000-4000 = $1000

1B.
Depreciation for 2017 = $55000 x 2/8 x 1/2 = $6875
Depreciation for 2018 = ($55000-6875) x 2/8 = $12031.25